OpenStudy (anonymous):

express using a positive exponent simplify the expression 8 negative 2power = evaluate

OpenStudy (saifoo.khan):

1/64

OpenStudy (anonymous):

saifoo.khan show me out to set this up for solving plse

OpenStudy (anonymous):

\[8^{-2}\] \[=1/8^{2}\] \[=1/64\]